A drunkard is walking along a straight road. he takes 6 steps forward and 4 steps backward and so on. each step is 1 m long and takes 1 s. there is apit 12 m away. the drunkard will fall into the pit after

A drunkard  is walking along a straight road. he takes 6 steps forward and 4 steps backward and so on. each step is 1 m long and takes 1 s. there is apit 12 m away. the drunkard will fall into the pit after

Grade:11

2 Answers

Arun
6 years ago
6 steps(F)- 4 steps(B) = 2 meters forward
2+ 6(F) -4 (B) = 4 meters forward
4+ 6(F) -4 (B) = 6 meters forward
6 + 6 (F) = 12 meters; fallen into a pit
Total steps = 36 steps
Hence after 36 seconds he will fall into a pit
Niraj Bishnoi
6 years ago
The drunkard will continue to move in the pattern 6steps forward 4steps backward, for the first 12-4=8m.
6-4 steps=2 steps = 2m which takes 6+2=8s.
If 2m takes 8s,
so, 8m will take 8s/2m x 8m= 8s x 4 = 32s.
He has travelled 8m, the remaining 4m will be travelled normally. SO that total time he takes becomes 32s + 4s= 36s